In a straight conductor of uniform crosss section charge q is flowing for time t. Let s be the specific charge of an electron. the momentum of all the free electrons per unit length of the conductor . due to their drift velocity only is 
 
Answer is q/ts

Grade:10

1 Answers

venketep prasad m c
7 years ago
@raghavendra  I = nAeVd  where n = no. of electrons per unit volume, A= area of cross section of the wire , e = charge of an electron , V= drift velocity
what we want is momentum per unit length of wire.(ie) mV, where m is the total mass of all electrons in the unit length of wire. here  I = (q/t) = nAeV. and m = n*(A*1)*(me). As n = no. of electrons per unit volume and A*1 is the volume of the unit length cross section and me is the mass of one electron.
Therefore mV = (nAVd)*m= (q*me)/(t*e)
As e/me = specific charge  = s,
mVd  = q/(t*s)